# Blocked Bloom Filters with Choices

### v1

@@ Abstract
We introduce Blocked Bloom filters with choices that work similarly to Blocked Bloom filters, except that for each key there are two (or more) alternative choices of blocks where the key’s information may be stored. The result is a filter that partially inherits the advantages of a Blocked Bloom filter, such as the ability to insert keys rapidly online or the ability to slightly overload the filter with only a small penalty to the false positive rate. At the same time, it avoids the major disadvantage of a Blocked Bloom filter, namely the larger space consumption. Our new data structure uses less space at the same false positive rate, or has a lower false positive rate at the same space consumption as a Blocked Bloom filter. We discuss the methodology, engineered implementation, a detailed performance evaluation and use cases in bioinformatics of Blocked Bloom filters with choices, showing that they can be of practical value.
@@ 1 Introduction
Bloom filters [2] are the most well known classical example of filters. Their overhead factor is relatively high with $C=1/\ln(2)\approx 1.443$. In addition, for $\varepsilon=2^{-k}$, Bloom filters access $k$ (deterministically computed) randomly distributed bits in memory to insert or query one key. Each of these accesses is likely a cache miss; so insertion and lookup speeds decrease with lower FPRs.
These disadvantages led to the development of other types of filters, including several variants of the Bloom filter, such as the Blocked Bloom filter [13], and filters based on different principles, such as the Cuckoo filter [6], (Vector) Quotient filter [1, 11], XOR filter [7], Binary Fuse filter [8], or Ribbon filter [5], to mention but a few. The goals were to reduce the space requirement (for the same FPR), or to improve insert or lookup speed, or all of these simultaneously. Another desirable property of a filter is extendability, i.e., its dynamic growth as more elements are inserted. The filters mentioned above are not extendable and have to be initialized with a given storage capacity, but more recent filters, such as Aleph filters [4] or Infini filters [3] are extendable, but come with different disadvantages.
@@
We leverage the “power of (two or more) choices” idea [10] to unite the advantages of Blocked and standard Bloom filters (fast speed, little extra overhead). The idea of two choices is used differently than for Cuckoo filters (which also use two choices to obtain better space efficiency at low FPRs): Our main idea is to use the structure of a Blocked Bloom filter, but choose between two blocks (given by two distinct block address hash functions) for each key. In this way, we obtain a more balanced block load, and we can optimize the placement of keys to re-use already used bits, together leading to a lower load overall and therefore a lower FPR. It will be shown in extensive evaluations that both effects together are strong enough to more than counteract the FPR doubling caused by searching at two different locations.
Article outline.
@@ 2.1 Bloom filters
Solving $p_{0}=1/2$ for $m$ yields $m=nk/\ln(2)\approx 1.443\,nk$; hence the well-known space overhead factor of $C\approx 1.443$, independektly of $k$.
The assumption behind the calculations above is that the $k$ hash functions behave as if they select $k$ distinct but otherwise independent random bit addresses. In practice, one can simply choose $k$ random functions from a sufficiently large family of simple hash functions. In very rare cases (for large $k$), two (or more) bit addresses may coincide, such that we effectively set or query less than $k$ distinct bits for a very small number of keys; this effect is so small that it has no measurable influence on the FPR when $m$ is large. However, this consideration can be important for Blocked Bloom filters.
@@
showing the smooth dependence of the FPR $\varepsilon$ on the overloading factor $\gamma$. For example, for 10% overload ($\gamma=1.1$) and $k=10$, we obtain an FPR of roughly twice the desired one ($1.912\cdot 2^{-10}$ instead of $2^{-10}$). Conversely, when we query an underfull Bloom filter, the FPR is even lower than the desired one.
@@ 2.2 Blocked Bloom filters
The obvious advantage of a Blocked Bloom filter over a standard Bloom filter is the increased speed, especially for low FPRs (high values of $k$). However, there is also a substantial disadvantage: For the same size as a Bloom filter, the overall FPR of a Blocked Bloom filter is higher due to two reasons: First, the $M$ blocks have different loads. Blocks with only few remaining unset bits have a high local FPR, while blocks with many remaining unset bits have a low local FPR. On average, because of Eq. (2) and Jensen’s inequality, the blocks with high load increase the average FPR more significantly than blocks with low load decrease it. The additional FPR cost increases with $k$; see Section 4 (Figure 4) for concrete numbers. The effect can be compensated by using more bits (hash functions) per key and/or more blocks, thus increasing the space overhead (but maintaining the desired FPR). Second, when independent random hash functions are chosen, the probability of a bit address collision with $k$ out of 512 can be substantially high, such that in several cases, only $k^{\prime}<k$ bits are effectively queried, increasing the FPR from $p_{1}^{-k}$ to $p_{1}^{-k^{\prime}}$. This effect can be compensated by choosing the bit addresses not independently, but in a way that ensures distinct addresses (see Section 3.3), at the cost of higher insertion and lookup times.
@@ 3.1 Basic idea
The idea behind the BlowChoc filter is that it works like a Blocked Bloom filter, but one has two (or more) independent choices for the block. The bits for a key may be set in either block. When querying the key, both blocks are examined, and the answer True is returned if all queried bits are set in at least one of the two blocks. In contrast to Cuckoo and Quotient filters, data of inserted keys is never moved after insertion.
@@ 4 Evaluation
All evaluations were run on a PC workstation with an AMD Ryzen 9 7950X3D CPU with 16 cores and hyperthreading, 128 GB of DDR5 memory with 4800 MHz and CL40 and a Western Digital Black SN850X SSD. After comparing different cost functions for insertion (Appendix C), the $\mbox{cost}^{\text{\footnotesize EXP}}_{\beta}$ function with $\beta\approx(1+\sqrt{5})/2$ produced the best overall FPRs. We hence only consider this particular cost function in the evaluations.
@@ Appendix C Evaluation of Cost Functions
For all cost functions, values of $k$, choices and parameters, the distinct bit selection strategy has a smaller FPR compared to the random bit selection strategy, as expected. The optimal cost parameter ($\sigma$ for $\mbox{cost}^{\text{\footnotesize MIX}}$, $\mu$ for $\mbox{cost}^{\text{\footnotesize LA}}$, and $\beta$ for $\mbox{cost}^{\text{\footnotesize EXP}}$) depends on $k$ and the number of choices. Using the mixed or lookahead cost functions, we are not able to achieve the target FPR of the standard Bloom filter with any parameter values. The exponential cost function ($\mbox{cost}^{\text{\footnotesize EXP}}$) leads to the overall best FPRs and with a surprisingly consistently good parameter value of $\beta\approx 1.6$ or the golden ratio $\beta\approx(1+\sqrt{5})/2$ for all choices, bit selection strategies, and $k$. In the main evaluations, we only consider $\mbox{cost}^{\text{\footnotesize EXP}}_{\beta}$ with $\beta=(1+\sqrt{5})/2$.
@@ Appendix D Genome Data Sources and Statistics
The human Homo sapiens genome has 3.1 billion base pairs of DNA and 2.5 billion distinct 31-grams. We used the following reference provided by the T2T consortium: https://s3-us-west-2.amazonaws.com/human-pangenomics/T2T/CHM13/assemblies/analysis_set/chm13v2.0.fa.gz

### v2

@@ Abstract
We introduce Blocked Bloom filters with choices that work similarly to Blocked Bloom filters, except that for each key there are two (or more) alternative choices of blocks where the key’s information may be stored. When inserting a key, we select the block using a cost function which takes into account the current load and the additional number of bits to be set in the candidate blocks. The result is a filter that partially inherits the advantages of a Blocked Bloom filter, such as the ability to insert keys rapidly online or the ability to slightly overload the filter with only a small penalty to the false positive rate. At the same time, it avoids the major disadvantage of a Blocked Bloom filter, namely the larger space consumption. Our new data structure uses less space at the same false positive rate, or has a lower false positive rate at the same space consumption as a Blocked Bloom filter. We discuss the methodology, cost functions for block selection, engineered implementation, a detailed performance evaluation and use cases in bioinformatics of Blocked Bloom filters with choices, showing that they can be of practical value.
@@ 1 Introduction
Bloom filters [2] are the most well known classical example of filters. Their overhead factor is relatively high with $C=1/\ln(2)\approx 1.443$. In addition, for $\varepsilon=2^{-k}$, Bloom filters access $k$ (deterministically computed) randomly distributed bits in memory to insert or query one key. Each of these accesses is a likely cache miss; so insertion and lookup speeds decrease with lower FPRs.
These disadvantages led to the development of other types of filters, including several variants of the Bloom filter, such as the Blocked Bloom filter [16], and filters based on different principles, such as the Cuckoo filter [7], (Vector) Quotient filter [1, 14], XOR filter [8], Binary Fuse filter [9], or (Bumped) Ribbon filter [6, 5], to mention but a few. The goals were to reduce the space requirement (for the same FPR), or to improve insert or lookup speed, or all of these simultaneously. Another desirable property of a filter is extendability, i.e., its dynamic growth as more elements are inserted. The filters mentioned above are not extendable and have to be initialized with a given storage capacity, but more recent filters, such as Aleph filters [4] or Infini filters [3] are extendable, but come with different disadvantages.
@@
The power of choices.
We leverage the “power of (two or more) choices” idea [13] to unite the advantages of Blocked and standard Bloom filters (fast speed, little extra overhead). The idea to use choices within Bloom filters is not entirely new. It was already explored for standard (not Blocked) Bloom filters, using two independent sets of bit positions in the entire bit array [12]. This approach needs $2k$ memory accesses with cache misses instead of the already large $k$, further reducing throughput. The experimental results also showed no distinctive other advantages. In another approach, balanced Blocked Bloom filters [10], a key is placed in the first block which stores less than a pre-defined number of keys. A counter monitors the number of keys in each block. Counting keys is faster but less informative than looking at actual bit usage, and the approach led to only small improvements.
@@
Our main contribution is to use the structure of a Blocked Bloom filter, but choose between two blocks (given by two distinct block address hash functions) based on a new engineered cost function. In this way, we obtain a more balanced block load, and we can optimize the placement of keys to re-use already used bits, together leading to a lower load overall and therefore a lower FPR. It will be shown in extensive evaluations that the effects of optimizing the key placements using a cost function are strong enough to more than counteract the FPR doubling caused by searching at two different locations. This is in contrast to balanced Blocked Bloom filters, which do not optimize the key placement and therefore do not efficiently re-use already set bits, resulting in a filter with a higher FPR compared to Bloom filters [10].
@@ 2.1 Bloom filters
Solving $p_{0}=1/2$ for $m$ yields $m=nk/\ln(2)\approx 1.443\,nk$; hence the well-known space overhead factor of $C\approx 1.443$, independently of $k$.
The assumption behind the calculations above is that the $k$ hash functions behave as if they select $k$ distinct but otherwise independent random bit addresses. In practice, one can simply choose $k$ random functions from a sufficiently large family of simple hash functions. In very rare cases, two (or more) bit addresses may coincide, such that we effectively set or query less than $k$ distinct bits for a very small number of keys; this effect is so small that it has no measurable influence on the FPR when $m$ is large and $k$ is a small constant. However, this consideration can be important for Blocked Bloom filters.
@@
showing the smooth dependence of the FPR $\varepsilon$ on the overloading factor $\gamma$. For example, for 10% overload ($\gamma=1.1$) and $k=10$, we obtain an FPR of roughly twice the desired one ($1.912\cdot 2^{-10}$ instead of $2^{-10}$). Conversely, when we query an underfull Bloom filter, the FPR is better (lower) than the desired one.
@@ 2.2 Blocked Bloom filters
The obvious advantage of a Blocked Bloom filter over a standard Bloom filter is the increased speed, especially for low FPRs (high values of $k$). However, there is also a substantial disadvantage. For the same size as a Bloom filter, the overall FPR of a Blocked Bloom filter is higher due to two reasons: First, the $M$ blocks have different loads. Blocks with only few remaining unset bits have a high local FPR, while blocks with many remaining unset bits have a low local FPR. On average, because of Eq. (2) and Jensen’s inequality, the blocks with high load increase the average FPR more significantly than blocks with low load decrease it. The surplus FPR caused by inhomogeneity increases with $k$; see Section 4 (Figure 4) for concrete numbers. The effect can be compensated by using more bits (hash functions) per key and/or more blocks, thus increasing the space overhead (but maintaining the desired FPR). Second, when independent random hash functions are chosen, the probability of a bit address collision with $k$ out of 512 can be substantially high, such that in several cases, only $k^{\prime}<k$ bits are effectively queried, increasing the FPR from $p_{1}^{-k}$ to $p_{1}^{-k^{\prime}}$. This effect can be compensated by choosing the bit addresses not independently, but in a way that ensures distinct addresses (see Section 3.3), at the cost of higher insertion and lookup times.
@@ 3.1 Basic idea
The idea behind the BlowChoc filter is that it works like a Blocked Bloom filter, but there are two (or more) independent choices for the block. The bits for a key may be set in either block. When querying the key, both blocks are examined, and the answer True is returned if all queried bits are set in at least one of the blocks. In contrast to Cuckoo and Quotient filters, data of inserted keys is never moved after insertion.
@@ 4 Evaluation
All evaluations were run on a PC workstation with an AMD Ryzen 9 7950X3D CPU with 16 cores and hyperthreading, 128 GB of DDR5 memory with 4800 MHz and CL40 and a Western Digital Black SN850X SSD. After comparing different cost functions for insertion (Appendix C), the $\mbox{cost}^{\text{\footnotesize EXP}}_{\beta}$ function with $\beta\approx(1+\sqrt{5})/2$ produced the best overall FPRs. The appearance of the golden ratio is purely coincidental. We only consider this particular cost function in the evaluations.
@@ Appendix C Evaluation of Cost Functions
For all cost functions, values of $k$, choices and parameters, the distinct bit selection strategy has a smaller FPR compared to the random bit selection strategy, as expected. The optimal cost parameter ($\sigma$ for $\mbox{cost}^{\text{\footnotesize MIX}}$, $\mu$ for $\mbox{cost}^{\text{\footnotesize LA}}$, and $\beta$ for $\mbox{cost}^{\text{\footnotesize EXP}}$) depends on $k$ and the number of choices. Using the mixed or lookahead cost functions, we are not able to achieve the target FPR of the standard Bloom filter with any parameter values. The exponential cost function ($\mbox{cost}^{\text{\footnotesize EXP}}$) leads to the overall best FPRs and with a surprisingly consistently good parameter value of $\beta\approx 1.6$ or the golden ratio $\beta\approx(1+\sqrt{5})/2\approx 1.618$ for all choices, bit selection strategies, and $k$. In the main evaluations, we only consider $\mbox{cost}^{\text{\footnotesize EXP}}_{\beta}$ with $\beta=(1+\sqrt{5})/2$. The fact that close-to-optimal values for $\beta$ are close to the golden ratio appears to be purely coincidental.
@@ Appendix D Genome Data Sources and Statistics
The human Homo sapiens genome has 3.1 billion base pairs of DNA and 2.5 billion distinct 31-grams. We used the following reference provided by the T2T consortium [17]: https://s3-us-west-2.amazonaws.com/human-pangenomics/T2T/CHM13/assemblies/analysis_set/chm13v2.0.fa.gz